OpenStudy (anonymous):

b/b^2-25 + 5/b+5 - 6/b

OpenStudy (anonymous):

\[\frac{b}{b^2-25}+\frac{5}{b+5}-\frac{6}{b}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry i'm not sure how you to type it out like that

OpenStudy (anonymous):

alright. \[\frac{b(b+5) + 5(b-5)(b+5)}{(b+5)(b-5)(b+5)}-\frac{6}{b}\]

OpenStudy (anonymous):

\[\frac{b + 5b-25}{(b-5)(b+5)}-\frac{6}{b}\]

OpenStudy (anonymous):

\[=\frac{b(6b-25) - 6(b-5)(b+5)}{b(b-5)(b+5)}\]

OpenStudy (anonymous):

\[=\frac{6b^2-25b-6b^2+25}{b(b-5)(b+5)}\] \[=\frac{-25b+25}{b(b-5)(b+5)}\]

OpenStudy (anonymous):

but I think \[\frac{6b-25}{(b-5)(b+5)}-\frac{6}{b}\] is simple enough

OpenStudy (anonymous):

neither of those is an answer on the sheet

OpenStudy (anonymous):

what are the answers?

OpenStudy (anonymous):

are you sure the question is \[\frac{b}{b^2-25}+\frac{5}{b+5}-\frac{6}{b}\] because otherwise, the answer could be very different.

OpenStudy (anonymous):

6b^2-25b+150/b(b+5)(b-5) 25(b+6)/b(b+5)(b-5) 25(b-6)/(b+5)(b-5) -25(b-6)/b(b+5)(b-5)

OpenStudy (anonymous):

yes, that is the question

OpenStudy (anonymous):

oh those are good, because we're almost there, just one more step (factor out the 25) \[\frac{6b^2 -25b -6(b^2-25)}{b(b-5)(b+5)}\]\[=\frac{-25b + 25}{b(b-5)(b+5)}\]\[=\frac{-25(b - 1)}{b(b-5)(b+5)}\] I think the answer is this one: -25(b-6)/b(b+5)(b-5)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yeah I checked and the answer comes up as -25(b-6)/b(b+5)(b-5) (I did a mistake somewhere so the 6 ended up as a 1) :-(

OpenStudy (anonymous):

no worries, i'd give you 2345098 medals if i could